OpenStudy (anonymous):

Find the slope of the line that passes through the pair of points (17, 9) and (5, 29).

OpenStudy (nurali):

\[m=\frac{ y _{2}-y_{1} }{ x _{2}-x_{1} }\] (x1,y1) and (x2,y2) (17, 9) and (5, 29) Put in Formula \[m=\frac{ 29-9 }{ 5-17 }\] \[m=?\]

OpenStudy (akashdeepdeb):

(x1,y1) = (17,9) (x2,y2) = (5,29) Slope: (y2-y1)/(x2-x1) Understood? :)

OpenStudy (anonymous):

would you show me the answer

OpenStudy (akashdeepdeb):

(y2-y1)/(x2-x1) = (29-9)/(5-17) = 20/-12 = -5/3 Now get it? :)

OpenStudy (anonymous):

yes:)

OpenStudy (akashdeepdeb):

:)

OpenStudy (anonymous):

WHATS 10- -9 I AINT GOT CALCULATOR

OpenStudy (akashdeepdeb):

what?

OpenStudy (anonymous):

Find the slope of the line that passes through the pair of points (–1, –9) and (–3, 10).